# The Multi-Wavelength Context of Delayed Radio Emission in TDEs: Evidence for Accretion-Driven Outflows

### v1

@@ Appendix A VLASS search for late-rising radio emission from known TDEs
Figure 6: The ages of the TDEs in our VLASS sample at the time each was observed in VLASS epoch 1. We detect radio emission from 5 TDEs out of a total sample of 91. Three detections are likely host galaxy emission (NGC 5905, F01004-2237, and J100933), while two are confirmed to be transient radio emission associated with the TDE (Sw J1644+57 and ASASSN-15oi). Both confirmed TDE radio detections occur on timescales of $2-6$ years post-discovery (rest frame), suggesting that this may be an interesting timescale for deeper follow up. (A sixth TDE, AT2018hyz, is detected in VLASS epoch 3 at a similar timescale post-disruption.)
@@ Appendix B Data Tables
In this section, we present a table listing the original references for all published optical/UV, X-ray, and radio data used in our analysis (Table 4) and complete data tables for all new observations. Our new radio observations of the TDE AT2019qiz (Section 2.1) are listed in Table 5 and new Swift-XRT observations of our sample (Section 2.2) are listed in Table 6. Table 6 also includes public archival Swift observations of our TDEs that have not previously been published, which were reduced using the same method as our new observations (Section 2.2). We will also make the full multi-wavelength dataset for each TDE in our sample publicly available online through OTTER (Franz et al in prep).
@@
Our radio data first published in C24 were taken with the VLA (programs 20B-492, PI: Alexander; 21-303, PI: Hajela; 22B-205, PI: Cendes), MeerKAT (programs SCI-20210212-YC-01, DDT-20220414-YC-01, and SCI-20220822-YC-01; PI: Cendes), and the Australia Telescope Compact Array (ATCA; programs C3472, PI: Cendes and C3325, PI: Alexander). The full list of observations, measured radio flux densities and upper limits, and details of the data reduction for these 24 objects are given in Cendes et al. (2022b), C24, and references therein. Our new radio observations of AT2019qiz are described in Section 2.1; our measured flux densities and the observing program used to collect each data point are listed in Table 5.
@@
Sixteen of our TDEs were included in the uniform X-ray analysis published by Guolo et al. (2023), which includes observations taken by XMM-Newton and by the X-ray Telescope (XRT, Burrows et al. 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004). We use the data from Guolo et al. (2023) for 14 of these objects; for ASASSN-15oi, we instead use the more comprehensive dataset presented by Hajela et al. (2024), and for AT2019qiz, we use the X-ray dataset from Nicholl et al. (2024), which corrects the XRT photometry for the presence of a nearby background source unrelated to the TDE. We also include published observations from the Chandra X-ray Observatory for a few TDEs: ASASSN-14ae (Jonker et al., 2020), PS16dtm (Petrushevska et al., 2023), AT2017eqx (Nicholl et al., 2019), and AT2018hyz (Cendes et al., 2022b). For AT2018fyk, we also include the late-time X-ray dataset from Pasham et al. (2024) consisting of XRT, XMM, and Chandra observations at $t>1000$ d post-discovery. For ASASSN-19bt, we report XMM-Newton photometry from Holoien et al. (2019a) and Swift-XRT and Chandra photometry from Christy et al. (2024).
@@
We include all of the optical/UV photometry from Nicholl et al. (2022) for the TDEs presented in that work (Table 4 also lists the original references for these data). For AT2018lna, AT2019teq, AT2020mot, AT2020opy, AT2020pj, and AT2020wey we include $g$ and $r$-band photometry from the Zwicky Transient Facility (ZTF; Bellm et al. 2019b) obtained from the Automatic Learning for the Rapid Classification of Events (ALeRCE) broker (Förster et al., 2021) and UVOT photometry from Hammerstein et al. (2023a). For AT2020neh, AT2020nov, and AT2020vwl, we include only $g$ and $r$-band photometry from ZTF from ALeRCE. For OGLE17aaj we include I-band photometry from Optical Gravitational Lensing Experiment (OGLE; Wyrzykowski et al. 2014) up until MJD = 58200. We include the UVOT photometry from Wevers et al. (2019b) for AT2018fyk and from van Velzen et al. (2021) for AT2018bsi. For ASASSN-19bt, we include both published photometry and the public light curve from the ASAS-SN Sky Patrol (Hart et al., 2023). Six TDEs in our sample (ASASSN-14li, ASASSN-15oi, AT2018hyz, ASASSN-19bt, AT2019teq, and AT2019ehz) show late-time plateaus in their optical and UV light curves that are not well-captured by MOSFiT, as discussed in Section 3.2. We therefore only use the early photometry from these events for our modeling. The cutoff date for each of these TDEs is listed in Table 4. Similarly, for PS16dtm we include only the photometry from Blanchard et al. (2017) as the additional photometry published by Petrushevska et al. (2023) is not well-captured by MOSFiT.
@@
Table 4: References for all data used in our analysis. All data for which the reference is “This work” are given in Tables 5 and 6.
@@
AT2018zr van Velzen et al. (2019b); Guolo et al. (2023) van Velzen et al. (2019b);
Holoien et al. (2019b); C24
@@
AT2018fyk Wevers et al. (2019b) Guolo et al. (2023); Wevers et al. (2019b);
@@
AT2018hyz Gomez et al. (2020); Cendes et al. (2022b); Horesh et al. (2018);
@@
Hammerstein et al. (2023a)
ASASSN-19bt ASAS-SN Sky Patrol; Holoien et al. (2019a); Christy et al. (2024)
Holoien et al. (2019a); Christy et al. (2024)
@@
AT2019teq ZTF $g,r$ (ALERCE); Guolo et al. (2023); Cendes et al. (2022a);
Hammerstein et al. (2023a); This work C24
@@
Hammerstein et al. (2023a) This work
@@
Hammerstein et al. (2023a) This work
@@
Table 6: New Swift-XRT data used in this work. This table includes both our new late-time TOO observations and public archival observations re-reduced by us. The date column gives the mean observation time for each data point, while the time since discovery includes the observation length or the full time range for stacked observations. The luminosity is calculated assuming a $\Gamma=2$ power-law spectrum, as discussed in Section 2.2. All upper limits are $3\sigma$ (Gaussian-equivalent).

### v2

@@ Appendix A VLASS search for late-rising radio emission from known TDEs
Figure 6: The ages of the TDEs in our VLASS sample at the time each was observed in VLASS epoch 1. We detect radio emission from 5 TDEs out of a total sample of 91. Three detections are likely host galaxy emission (NGC 5905, F01004-2237, and J100933), while two are confirmed to be transient radio emission associated with the TDE (Sw J1644+57 and ASASSN-15oi). Both confirmed TDE radio detections occur on timescales of 262-6 years post-discovery (rest frame), suggesting that this may be an interesting timescale for deeper follow up. (A sixth TDE, AT2018hyz, is detected in VLASS epoch 3 at a similar timescale post-disruption.)
@@ Appendix B Data Tables
In this section, we present a table listing the original references for all published optical/UV, X-ray, and radio data used in our analysis (Table 6) and complete data tables for all new observations. Our new radio observations of the TDE AT2019qiz (Section II.1) are listed in Table 7 and new Swift-XRT observations of our sample (Section II.2) are listed in Table 8. Table 8 also includes public archival Swift observations of our TDEs that have not previously been published, which were reduced using the same method as our new observations (Section II.2). We will also make the full multi-wavelength dataset for each TDE in our sample publicly available online through OTTER (the Open mulTiwavelength Transient Event Repository; Franz et al. 2025). OTTER has both a web application and a Python API for easy data access.
@@
The radio data first published in C24 were taken with the VLA (programs 20B-492, PI: Alexander; 21-303, PI: Hajela; 22B-205, PI: Cendes), MeerKAT (programs SCI-20210212-YC-01, DDT-20220414-YC-01, and SCI-20220822-YC-01; PI: Cendes), and the Australia Telescope Compact Array (ATCA; programs C3472, PI: Cendes and C3325, PI: Alexander). The full list of observations, measured radio flux densities and upper limits, and details of the data reduction for these 24 objects are given in Cendes et al. (2022a), C24, and references therein. Our new radio observations of AT2019qiz are described in Section II.1; our measured flux densities and the observing program used to collect each data point are listed in Table 7.
@@
Sixteen of our TDEs were included in the uniform X-ray analysis published by Guolo et al. (2024), which includes observations taken by XMM-Newton and by the X-ray Telescope (XRT, Burrows et al. 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004). We use the data from Guolo et al. (2024) for 14 of these objects; for ASASSN-15oi, we instead use the more comprehensive dataset presented by Hajela et al. (2025), and for AT2019qiz, we use the X-ray dataset from Nicholl et al. (2024), which corrects the XRT photometry for the presence of a nearby background source unrelated to the TDE. We also include published observations from the Chandra X-ray Observatory for a few TDEs: ASASSN-14ae (Jonker et al., 2020), PS16dtm (Petrushevska et al., 2023), AT2017eqx (Nicholl et al., 2019), and AT2018hyz (Cendes et al., 2022a). For AT2018fyk, we also include the late-time X-ray dataset from Pasham et al. (2024) consisting of XRT, XMM, and Chandra observations at $t>1000$ d post-discovery. For ASASSN-19bt, we report XMM-Newton photometry from Holoien et al. (2019b) and Swift-XRT and Chandra photometry from Christy et al. (2024).
@@
We include all of the optical/UV photometry from Nicholl et al. (2022) for the TDEs presented in that work (Table 6 also lists the original references for these data). For AT2018lna, AT2019teq, AT2020mot, AT2020opy, AT2020pj, and AT2020wey we include $g$ and $r$-band photometry from the Zwicky Transient Facility (ZTF; Bellm et al. 2019a) obtained from the Automatic Learning for the Rapid Classification of Events (ALeRCE) broker (Förster et al., 2021) and UVOT photometry from Hammerstein et al. (2023b). For AT2020neh, AT2020nov, and AT2020vwl, we include only $g$ and $r$-band photometry from ZTF from ALeRCE. For OGLE17aaj we include I-band photometry from Optical Gravitational Lensing Experiment (OGLE; Wyrzykowski et al. 2014) up until MJD = 58200. We include the UVOT photometry from Wevers et al. (2019a) for AT2018fyk and from van Velzen et al. (2021) for AT2018bsi. For ASASSN-19bt, we include both published photometry and the public light curve from the ASAS-SN Sky Patrol (Hart et al., 2023). Six TDEs in our sample (ASASSN-14li, ASASSN-15oi, AT2018hyz, ASASSN-19bt, AT2019teq, and AT2019ehz) show late-time plateaus in their optical and UV light curves that are not well-captured by MOSFiT, as discussed in Section III.2. We therefore only use the early photometry from these events for our modeling. The cutoff date for each of these TDEs is listed in Table 6. Similarly, for PS16dtm we include only the photometry from Blanchard et al. (2017) as the additional photometry published by Petrushevska et al. (2023) is not well-captured by MOSFiT.
@@
Table 6: References for all data used in our analysis. All data for which the reference is “This work” are given in Tables 7 and 8.
@@
AT2018zr van Velzen et al. (2019a); Guolo et al. (2024) van Velzen et al. (2019a);
Holoien et al. (2019a); C24
@@
AT2018fyk Wevers et al. (2019a) Guolo et al. (2024); Wevers et al. (2019a);
@@
AT2018hyz Gomez et al. (2020); Cendes et al. (2022a); Horesh et al. (2018);
@@
Hammerstein et al. (2023b)
ASASSN-19bt ASAS-SN Sky Patrol; Holoien et al. (2019b); Christy et al. (2024)
Holoien et al. (2019b); Christy et al. (2024)
@@
AT2019teq ZTF $g,r$ (ALERCE); Guolo et al. (2024); Cendes et al. (2022b);
Hammerstein et al. (2023b); This work C24
@@
Hammerstein et al. (2023b) This work
@@
Hammerstein et al. (2023b) This work
@@
Table 8: New Swift-XRT data used in this work. This table includes both our new late-time TOO observations and public archival observations re-reduced by us. The date column gives the mean observation time for each data point, while the uncertainty on the time since discovery is taken to be either the observation length for a single observation or the full time range for stacked observations. The luminosity is calculated assuming a $\Gamma=2$ power-law spectrum, as discussed in Section II.2. All upper limits are $3\sigma$ (Gaussian-equivalent). This is a subset of the full table, which will be available online in machine-readable format.